A 25-g copper penny is given a charge of-4.0 penny? 10-9 C. How many excess electrons are on the x 9.426 x 1010 electrons 4.296 x 1010 clectrons 2. 496 x 1010 electrons 1.946 x 1010 electrons Question 2 (12.5 points) atomic nucleus are typically 10-15 m apart. What is the electric force of repulsion between Protons in a nuclear protons? 320.7 N 430.7 N 230.7 N 130.7 N
Answer #1

1)

charge Q =N*e

N =Q/e

=(-4*10^-9 C)/(-1.602*10^-19 C/electron)

=2.496*10^10

option(C) is correct

2)

From coulombs law

F =k*q1*q2/(r^2)

=(8.99*10^9 N.m^2/C^2)*((1.602*10^-19 C)/(10^-15 m))^2

= 230.7 N

option(C) is correct